Your company is considering an investment that will have the following yearend cash flows: year 1, $8000; year 2, $10,000; year 3, 12,000. The company wants to earn a 15% annual return on its investment. How much should it pay for the investment, assuming quarterly compounding?
Answer #1

We use the formula:
A=P(1+r/4)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

8000=P1*(1+0.15/4)^(4*1)

P1=8000/(1+0.15/4)^(4*1)

=8000*0.863073095

=6904.58(Approx)

10,000=P2*(1+0.15/4)^(4*2)

P2=10,000/(1+0.15/4)^(4*2)

=$10,000*0.744895167

=$7448.95(Approx)

12000=P3(1+0.15/4)^(4*3)

P3=12000/(1+0.15/4)^(4*3)

=$12000*0.642898978

=$7714.79(Approx).

Hence total present value =6904.58+7448.95+$7714.79

=$22068.32(Approx).